OpenStudy (anonymous):

A funny artifact of complex analysis.\[\sum_{k=1}^nk=\frac{n(n+1)}2\]Which naturally goes to infinity, but via zeta function summation and Ramanujan sums, you get a value of \(-\frac1{12}\).

OpenStudy (anonymous):

1+2+3+4+...=-1/12 So that is TRUE?

OpenStudy (anonymous):

That's what Euler thought. :) http://en.wikipedia.org/wiki/Zeta_function_regularization#Definition

OpenStudy (anonymous):

http://plato.stanford.edu/entries/goedel/#IncThe Math is funny.

OpenStudy (anonymous):

So, at infinity number starts to act differently?

OpenStudy (experimentx):

pretty soon I need to work on these things.

OpenStudy (anonymous):

My condolences, haha. Unless you enjoy particle physics. If you do, ignore me.

OpenStudy (experimentx):

No I enjoy mathematical physics.

OpenStudy (experimentx):

you math major @badreferences ?

OpenStudy (anonymous):

I'm looking to do statistical physics in graduate school.

OpenStudy (experimentx):

so you are also physics major!! didn't know that!!

OpenStudy (anonymous):

:D

OpenStudy (experimentx):

I was looking for solution for "Topics in Algebra" by IN Herstein.

OpenStudy (anonymous):

Algebra is not my strong suit. I went for the analysis and stochastics route in undergraduate.

OpenStudy (experimentx):

my undergrad was mess!! still i need to do few things in math before tackle physics. It seems i still lack strong mathematical background for tackling physics.

OpenStudy (anonymous):

If I can recommend two introductory texts at graduate level, "All of Statistics" ( http://www.amazon.com/All-Statistics-Statistical-Inference-Springer/dp/0387402721 ) and "Mathematics: Its Content, Methods and Meaning" ( http://www.amazon.com/Mathematics-Content-Methods-Meaning-Dover/dp/0486409163/ref=sr_1_1?s=books&ie=UTF8&qid=1349543762&sr=1-1&keywords=mathematics+its+content+methods+and+meaning ). They should provide sufficient math background, I think. But I'm no expert on the subject.

OpenStudy (anonymous):

Shhh... I have illegal electronic copies.

OpenStudy (anonymous):

And the real things as well, so I don't feel to bad. haha.

OpenStudy (experimentx):

haha ... no probs!! i also own illegal electronic books!! I didn't take stat seriously ... I need to revise statistics and numerical analysis again.

OpenStudy (experimentx):

I need to revise whole physics i read in first year. But I've math exam first. Analysis II and Algebra II I have Tom Aposle's Mathematical analysis for Analysis .. everytime i try to read it ... i feels like reading "War and Peace"

OpenStudy (experimentx):

Same goes for Algebra, I have Linear Algebra by Serge Lang and Topics in algebra by IN Herstein ... but undergraduate level. But I'm doing Linear Algebra from MIT OCW.

OpenStudy (anonymous):

Hmm, the best analysis texts are Spivak's "Calculus" and "Calculus of Manifolds".

OpenStudy (anonymous):

I don't know about linear algebra. The best review of physics I can think of are Feynman's lectures and Dirac's books.

OpenStudy (experimentx):

Looks like I need to hunt down those books!!

OpenStudy (anonymous):

Wait a second. And hush, don't tell anyone. I'll post them here.

OpenStudy (experimentx):

thanks @badreferences for goodreferences

OpenStudy (anonymous):

Then what is: 1*2*3*4*5...

OpenStudy (experimentx):

lol ... that diverges. same goes from 1 + 2 + 3 + ... but if you view it as generating function of 1/(1+x)^2 = 1 + 2x + 3x^2 + ... it seems to converge ... there was some Question posten my mikael on this.

OpenStudy (anonymous):

Ya....... I remember it.

OpenStudy (experimentx):

I haven't encountered Riemann zeta function on my studies except for finding the particular values line zeta(2), zeta(4) while integrating some functions. I think it's better to go step by step on this.

OpenStudy (experimentx):

things aren't much intuitive ... there seems to lots of odd behaviours with this function which i don't understand. If you explain it ... probably you will be greatest mathematician of this time http://en.wikipedia.org/wiki/Millennium_Prize_Problems

OpenStudy (experimentx):

probably long way through analysis and algebra just to understand these stuff.

OpenStudy (experimentx):

Woops!! @mahmit2012 is here ...

OpenStudy (anonymous):

hello everybody.

OpenStudy (experimentx):

hello

OpenStudy (experimentx):

where do i start with Riemann zeta function?

OpenStudy (anonymous):

it is very simple because the series is divergence.

OpenStudy (anonymous):